Find the inverse of the function.
g(x)=-1/4x-9/5

Answer by CubeyThePenguin(3113)

y = -x/4 - 9/5
y + 9/5 = -x/4
-4(y + 9/5) = x

f^(-1)(x) = -4x - 36/5